Joe mixes 14 oz. of peanuts at $1.20 per oz with some raisins at $0.7 per oz. to get a mixture at $1.05 per oz. How many ounces of raisins are put in the mixture?
Answer by Boreal(15235)

14(1.20)+x(0.70)=(14+x)(1.05); the weight times the price has to be equal.
16.8+0.7 x=14.70+1.05 x
2.10=0.35x
x=6 oz.
6 oz*0.70 per ounce is $4.20
14 oz @1.20 is $16.80
That is 20 ounces for $21.00, or $1.05 per ounce